A. On May 1, 2018, Love Corporation declared a $58,600 cash dividend to be paid on May 31 to shareholders of record on May 15.

Required:

Record the events occurring on May 1 and May 31 in a horizontal statements model. In the Cash Flow column, indicate whether the item is an operating activity (OA), investing activity (IA), or financing activity (FA). If an element was not affected by the event, leave the cell blank. (Enter any decreases to account balances with a minus sign.)

LOVE CORPORATION
Horizontal Statements Model
Balance Sheet Income Statement Statement of Cash Flow
Date Assets = Liabilities + Common Stock + Retained Earnings Revenue Expense = Net Income
May 1 = + + =
May 15 = + + =
May 31 = + + =

B. Adams Company, which sells electric razors, had $350,000 of cost of goods sold during the month of June. The company projects a 9 percent increase in cost of goods sold during July. The inventory balance as of June 30 is $18,000, and the desired ending inventory balance for July is $19,000. Adams pays cash to settle 80 percent of its purchases on account during the month of purchase and pays the remaining 20 percent in the month following the purchase. The accounts payable balance as of June 30 was $38,000.

Required

  1. Determine the amount of purchases budgeted for July.

  2. Determine the amount of cash payments budgeted for inventory purchases in July.



a. Budgeted purchases

b. Cash payments
Answer #1

Solution A:

LOVE CORPORATION
Horizontal Statements Model
Balance Sheet Income Statement Statement of Cash Flow
Date Assets = Liabilities + Common Stock + Retained Earnings Revenue Expense = Net Income
May 1 = 58600 + + - 58600 =
May 15 = + + =
May 31 -58600 = - 58600 + + -58600 FA

Solution B:

Budgeted purchases = July cost of goods sold + ending inventory- beginning inventory

= $350000*109% + $19000 - $18000 = $382,500

Cash payments = accounts payable + July purchase *80% = $38000 + $382500*80% = $344,000